Projects/ Work Flow
- total Syk protocol now in place -- problems?
- supernatants being accumulated -- long term storage?
- fixed cells for Treg -- Kenny working through; will need to plan for analysis
- ICATA/ BioCSI/ URECA
- Ag uptake analysis due
- status of BioCSI basophil data?
- Shashi to review with IRB
- Neisha will redirect to Arah2 collaboration
- Lara out any day -- are we ready?
- contamination issues?
- ITN -- status? involvement from Julie?
Recruitment/ Team Update
- Performance reviews over next week
- review procedure for HTS installation
- discussion of centralized flow staff
- pipette calibration
Previous ToDo Items
- create protocol for total Syk measurement (wayne) -- done
- Re-organize data between the three Macs and add more back ups
New ToDo Items